If the diffusion coefficient of oxygen in the vitreous humour in the eye is 4.2×10−7 m2 s−1 how long does it take oxygen to diffuse an average distance of 1.0 mm through the vitreous humour? (to 2 s.f and in s)
1
Expert's answer
2020-04-24T10:51:14-0400
"T=\frac{x^2}{2D}"

"T=\frac{0.001^2}{2(4.2\cdot 10^{-7})}=1.2\ s"